OpenStudy (anonymous):

select the expressions that are equivalent to 7x^0. select all that apply. a. 7(7x)^0 b. 1 c. 7 d. (7x)^0

OpenStudy (anonymous):

I think it is b and d

OpenStudy (anonymous):

C

OpenStudy (anonymous):

anything to the ^0 = 1

OpenStudy (anonymous):

So 7*x^0 = 7*1 = 7

OpenStudy (anonymous):

Okay so then does the problem d equals one to because that is all to the ^0

OpenStudy (anonymous):

No d = 1 because it's (7x)^0 not 7(x)^0 A is also right.

OpenStudy (anonymous):

(7x)^0 = 1

OpenStudy (anonymous):

Oh that makes sense thanks

OpenStudy (anonymous):

Np